OpenStudy (anonymous):

Prove the identity 1-tan(x)tan(y)) = (cos(x+y)/cos(x)cos(y)

OpenStudy (jamesj):

What is cos(x+y) equal to? I.e., expand it out.

OpenStudy (anonymous):

Using the formula cos(A+B)=cos A cos B - sin A sin B cos (x + y) cos(x)cos(y) - sin(x)sin(y) cos(x)cos(y) sin(x)sin(y) RHS = ---------- = ---------------------- = ------------ - ------------ cos(x)cos(y) cos(x)cos(y) cos(x)cos(y) cos(x)cos(y) sin(x) sin(y) = 1 - ------- * ------ = 1 - tan(x)*tan(y) = 1 - tan(x)tan(y) = LHS cos(x) cos(y)